# Heights and periods of algebraic cycles in families

### v1

@@ 1.2.1. Betti strata
(1.1) $S_{\mathscr{F}}(t):=\left\{s\in S({\mathbb{C}}):\dim_{\nu(s)}(\nu(S)\cap{\mathscr{F}}_{\mathrm{Betti}})\geq t\right\}$
@@
(1.2) $\emptyset=S_{{\mathscr{F}}}(\dim S+1)\subseteq S_{{\mathscr{F}}}(\dim S)\subseteq\cdots\subseteq S_{{\mathscr{F}}}(1)\subseteq S_{{\mathscr{F}}}(0)=S.$
@@
For each $t\geq 0$, $S_{\mathscr{F}}(t)$ is Zariski closed in $S$.
@@
In the Betti strata, $S_{{\mathscr{F}}}(1)$ plays a particularly important role. For example, by definition of the Betti foliation, $S_{{\mathscr{F}}}(1)$ contains any analytic curve $C\subseteq S^{\mathrm{an}}$ such that $\nu(C)$ is torsion. In particular, there are at most countably many $s\in S({\mathbb{C}})$ outside $S_{{\mathscr{F}}}(1)$ such that $\nu(s)$ is torsion.
Finally, let us point out that $S_{{\mathscr{F}}}(t)$ is closely related to the degeneracy loci defined by the first-named author. Indeed, when ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ is an abelian scheme, then $S_{{\mathscr{F}}}(t)$ is the $\nu(S)^{\mathrm{deg}}(-t+1)$ in [Gao20a, Defn. 1.6] for each $t\geq 1$.
@@ 1.2.2. Betti rank
A trivial upper bound for $r(\nu)$ is $r(\nu)\leq\min\{\dim S,\frac{1}{2}\dim{\mathbb{V}}_{{\mathbb{Q}},s}\}$ for any $s\in S({\mathbb{C}})$. One can also easily improve this trivial upper bound as will be explained below.
The Betti rank is easily seen to be related to the Betti strata in the following way: $\dim S-r(\nu)$ is the maximum of $t\geq 0$ such that $S_{{\mathscr{F}}}(t)$ contains a non-empty open subset of $S^{\mathrm{an}}$. Thus Theorem 1.4 implies that $r(\nu)=\dim S-\min\{t\geq 0:S_{{\mathscr{F}}}(t)=S\}$. However, this equality is, in general, not applicable to compute $r(\nu)$. But at this stage, we have
(1.5) $S_{{\mathscr{F}}}(1)\not=S\Leftrightarrow r(\nu)=\dim S.$
The second main theorem for the geometric part is the following formula for $r(\nu)$, which is often computable in practice. The VMHS ${\mathbb{E}}_{\nu}$ on $S$ induces a period map $\varphi=\varphi_{\nu}\colon S{\longrightarrow}\Gamma\backslash{\mathscr{D}}$, with ${\mathscr{D}}$ a mixed Mumford–Tate domain with Mumford–Tate group $\mathbf{G}$ (so ${\mathscr{D}}$ is a $\mathbf{G}({\mathbb{R}})^{+}$-orbit for $\mathbf{G}$ the generic Mumford–Tate group of the VMHS ${\mathbb{E}}_{\nu}$); we refer to §$\S$2.12.2 for the construction of $\varphi$ and ${\mathscr{D}}$.
Denote by $V$ the unipotent radical of $\mathbf{G}$. It equals ${\mathbb{V}}^{\prime}_{{\mathbb{Q}},s}$ for any $s\in S({\mathbb{C}})$, where ${\mathbb{V}}^{\prime}_{{\mathbb{Z}}}$ is the largest sub-VHS of ${\mathbb{V}}_{{\mathbb{Z}}}$ such that $\nu$ becomes torsion under the projection ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}/{\mathbb{V}}^{\prime}_{{\mathbb{Z}}})$; see Remark 2.2. Now the trivial upper bound on $r(\nu)$ can be easily improved to be $r(\nu)\leq\min\{\dim\varphi(S),\frac{1}{2}\dim V\}$.
Here is our formula to compute $r(\nu)$.
@@
Notice that the trivial upper bound on $r(\nu)$ above is recovered by taking $N=\{1\}$ and $N=\mathbf{G}$.
Here are two applications of Theorem 1.6, on two cases where the trivial upper bound on $r(\nu)$ is attained.
@@
Notice that in the situation of Corollary 1.7, we have $V={\mathbb{V}}_{{\mathbb{Q}},s}$.
Corollary 1.8 (Theorem 5.2).
Assume: (i) the connected algebraic monodromy group $H$ of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ is simple; (ii) $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ has no isotrivial sub-VHS, i.e. locally constant VHS. Then $r(\nu)=\min\{\dim\varphi(S),\frac{1}{2}\dim_{{\mathbb{Q}}}V\}$.
@@
Under the assumptions of either Corollary 1.7 or Corollary 1.8. Assume that $\varphi$ is generically finite and that $\dim S\leq\frac{1}{2}\dim_{{\mathbb{Q}}}V$. Then there exists a Zariski open dense subset $U$ of $S$ such that $\nu(s)$ is torsion for at most countably many $s\in U({\mathbb{C}})$. Indeed, one can take $U=S\setminus S_{{\mathscr{F}}}(1)$.
@@
If $\varphi$ is not necessarily generically finite, then the conclusion becomes: there exists a Zariski open dense subset $U$ of $S$ such that $\{s\in U({\mathbb{C}}):\nu(s)\text{ is torsion}\}$ is contained in at most countably many fibers of $\varphi$. Indeed, $\varphi(S)$ is an algebraic variety by [BBT23]. Then we can conclude by applying Corollary 1.9 to $\varphi(S)$ and the induced normal function $\nu^{\prime}$ (see §$\S$3.1 for the construction of $\nu^{\prime}$).
Applying either Corollary 1.7 or Corollary 1.8 to the Gross–Schoen and the Ceresa normal functions, we obtain the following corollary will be used to prove Theorem 1.1, Corollary 1.2 and Theorem 1.3.
Corollary 1.11 (Corollary 5.3).
If $g\geq 3$, then the Betti rank of the Gross-Schoen (resp. of the Ceresa) normal function is $33g-$ 3.
Hain [Ha24] gave a different proof of Corollary 1.11. When $g=3$, he even gave an explicit description on the locus where $\dim\nu_{\mathrm{Betti},s}(T_{s}S)<3g-3$.
@@ 1.3. Plan and Ingredients of proofs
The arithmetic part requires the construction of a suitable adelic line bundle $\overline{{\mathscr{L}}}$ over the base, a theory initiated by the second-named author and developed in further joint work with Yuan [YZ21], such that the height function $h_{\overline{{\mathscr{L}}}}$ is the Beilinsin–Bloch height. Then, the desired height inequalities will follow from suitable bigness properties of $\overline{{\mathscr{L}}}$. In the context of Theorem 1.1 and Corollary 1.2, by the second-named author’s [Zha10, Thm. 2.5.5] the desired $\overline{{\mathscr{L}}}$ can be obtained from a suitable Deligne pairing. The required bigness property is the bigness of the generic fiber $\tilde{{\mathscr{L}}}$ of $\overline{{\mathscr{L}}}$. We furthermore show in Proposition 6.3 that the bigness of $\tilde{{\mathscr{L}}}$ will follow from the non-vanishing of $c_{1}(\overline{{\mathscr{L}}})^{\wedge\dim S}$, with $c_{1}(\overline{{\mathscr{L}}})$ the curvature form. The arithmetic part of this paper is confined to the Gross-Shoen and the Ceresa cycles.
@@
The geometric part studies the (admissible) normal function $\nu$ associated with $Z$. The main results are the Zariski closedness of the Betti strata and the formula for the Betti rank $r(\nu)$. This part has been explained in §$\S$1.2. The proofs of Theorem 1.4 and Theorem 1.6 are simultaneous and follow the guideline of the first-named author’s [Gao20a] on the generic rank of the Betti map for abelian schemes. A core of our proof is Ax–Schanuel for VMHS independently proved by Chiu [Chi21] and Gao–Klingler [GK24], which will be used multiple times.
@@
Hain’s works on the Hodge-theoretic computation of the Archimedean local height pairing and the Betti form are the key ingredients to bridging the geometric and arithmetic parts. Indeed, Hain in [Hai90] proved that the archimedean local height pairing could be computed using the metricized biextension line bundle on $S$, and in [HR04] computed (joint with Reed), the curvature form of the metrized biextension line bundle. In our situation, we work with the height pairing of $Z_{s}$ with itself for any $s\in S(\overline{\mathbb{Q}})$, and the metrized biextension line bundle in question is the pullback of the metrized tautological line bundle under $\nu$. Hain in [Hai13] proved that its curvature form $\beta_{\nu}$, called the Betti form, is a semi-positive $(1,1)$-form. We show in Corollary D.7 that $\beta_{\nu}^{\wedge\dim S}\not\equiv 0$ if and only if $r(\nu)=\dim S$. Finally, back to the situation of Theorem 1.1, Corollary 1.2, and Theorem 1.3. Hain’s formula [Hai90, Prop. 3.3.12] and the second-named author’s [Zha10, Thm. 2.5.5] together imply that the curvature $c_{1}(\overline{{\mathscr{L}}})$ is precisely the Betti form. This finishes the bridge between the geometric part and the arithmetic part. Again, the bridge between the geometric part and the arithmetic part works in much larger generality, provided that the arithmetic part is solved.

### v2

@@ 1.2.1. Betti strata
(1.2) $S^{\mathrm{Betti}}(t):=\left\{s\in S({\mathbb{C}}):\dim_{\nu(s)}(\nu(S)\cap{\mathscr{F}}_{\mathrm{Betti}})\geq t\right\}$
@@
(1.3) $\emptyset=S^{\mathrm{Betti}}(\dim S+1)\subseteq S^{\mathrm{Betti}}(\dim S)\subseteq\cdots\subseteq S^{\mathrm{Betti}}(1)\subseteq S^{\mathrm{Betti}}(0)=S.$
@@
For each $t\geq 0$, $S^{\mathrm{Betti}}(t)$ is Zariski closed in $S$.
@@
In the Betti strata, $S^{\mathrm{Betti}}(1)$ plays a particularly important role. For example, $S^{\mathrm{Betti}}(1)$ contains any complex analytic curve $C\subseteq S$ such that $\nu(C)$ is torsion because all torsion multi-sections are leaves of the Betti foliation. In particular, there are at most countably many $s\in S({\mathbb{C}})$ outside $S^{\mathrm{Betti}}(1)$ such that $\nu(s)$ is torsion.
Finally, let us point out that $S^{\mathrm{Betti}}(t)$ is closely related to the degeneracy loci defined by the first-named author. Indeed, when ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ is an abelian scheme, then $S^{\mathrm{Betti}}(t)$ is the $\nu(S)^{\mathrm{deg}}(-t+1)$ in [Gao20a, Defn. 1.6] for each $t\geq 1$.
@@ 1.2.2. Betti rank
A trivial upper bound is $r(\nu)\leq\min\{\dim S,\frac{1}{2}\dim{\mathbb{V}}_{{\mathbb{Q}},s}\}$ for any $s\in S({\mathbb{C}})$. One can also easily improve this trivial upper bound, as explained below.
The Betti rank is easily seen to be related to the Betti strata in the following way: $\dim S-r(\nu)$ is the maximum of $t\geq 0$ such that $S^{\mathrm{Betti}}(t)$ contains a non-empty open subset of $S^{\mathrm{an}}$. Thus Theorem 1.4 implies that $r(\nu)=\dim S-\min\{t\geq 0:S^{\mathrm{Betti}}(t)=S\}$. However, this equality is, in general, not applicable to compute $r(\nu)$. But at this stage, we have
(1.6) $S^{\mathrm{Betti}}(1)\not=S\Longleftrightarrow r(\nu)=\dim S.$
The second main theorem for the geometric part is the following formula for $r(\nu)$, which is often computable in practice. The VMHS ${\mathbb{E}}_{\nu}$ on $S$ induces a period map $\varphi=\varphi_{\nu}\colon S{\longrightarrow}\Gamma\backslash{\mathscr{D}}$, with ${\mathscr{D}}$ a mixed Mumford–Tate domain whose Mumford–Tate group we denote by $\mathbf{G}$ (so ${\mathscr{D}}$ is a $\mathbf{G}({\mathbb{R}})^{+}$-orbit and $\mathbf{G}$ is the generic Mumford–Tate group of the VMHS ${\mathbb{E}}_{\nu}$); we refer to §$\S$2.12.2 for the construction of $\varphi$ and ${\mathscr{D}}$.
Denote by $V$ the unipotent radical of $\mathbf{G}$. It equals ${\mathbb{V}}^{\prime}_{{\mathbb{Q}},s}$ for any $s\in S({\mathbb{C}})$, where
[…]
See Remark 2.2. Now the trivial upper bound on $r(\nu)$ can be easily improved to be $r(\nu)\leq\min\{\dim\varphi(S),\frac{1}{2}\dim V\}$.
Here is our formula to compute $r(\nu)$. The trivial upper bound on $r(\nu)$ above is recovered by taking $N=\{1\}$ and $N=\mathbf{G}$.
@@
We will give two applications of Theorem 1.6 on two cases where the trivial upper bound on $r(\nu)$ is attained: Theorem 5.1 when $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ is irreducible, and Theorem 5.2 when the algebraic monodromy group of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ is simple (or when the Mumford–Tate group of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ is simple).
To highlight these applications, we state Theorem 5.1 in the introduction. We also point out that Theorem 5.2 could be more applicable in geometric situations.
@@
It is easy to see that (i) and (ii) of Corollary 1.7 imply ${\mathbb{V}}^{\prime}_{{\mathbb{Z}}}={\mathbb{V}}_{{\mathbb{Z}}}$, and hence $V={\mathbb{V}}_{{\mathbb{Q}},s}$. Moreover, if $r(\nu)=\frac{1}{2}\dim_{{\mathbb{Q}}}{\mathbb{V}}_{{\mathbb{Q}},s}$, then ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ must be an abelian scheme by Griffiths’ transversality; see the comment below (2.8).
Applying Corollary 1.7 to the Gross–Schoen and the Ceresa normal functions, we obtain the following corollary will be used to prove Theorem 1.1, Corollary 1.2 and Theorem 1.3.
Corollary 1.8 (Corollary 5.3).
If $g\geq 3$, then the Betti rank of the Gross-Schoen (resp. of the Ceresa) normal function over ${\mathscr{M}}_{g}$ is $33g-$ 3.
Hain [Hai24] gave a different proof of Corollary 1.8. Our Theorem 1.6 can also be applied to any subvariety $S\subseteq{\mathscr{M}}_{g}$.
1.2.3. Consequence on torsion loci
@@
Under the assumptions of either Corollary 1.7 or Theorem 5.2. Assume that $\varphi$ is generically finite and that $\dim S\leq\frac{1}{2}\dim_{{\mathbb{Q}}}V$. Then there exists a Zariski open dense subset $U$ of $S$ such that $\nu(s)$ is torsion for at most countably many $s\in U({\mathbb{C}})$. Indeed, one can take $U=S\setminus S^{\mathrm{Betti}}(1)$.
@@
In general, $\varphi$ is not necessarily generically finite. Then the conclusion becomes: there exists a Zariski open dense subset $U$ of $S$ such that $\{s\in U({\mathbb{C}}):\nu(s)\text{ is torsion}\}$ is contained in at most countably many fibers of $\varphi$. Indeed, $\varphi(S)$ is an algebraic variety by [BBT23]. Then we can conclude by applying Corollary 1.9 to $\varphi(S)$ and the induced normal function $\nu^{\prime}$ (see §$\S$3.1 for the construction of $\nu^{\prime}$).
In Corollary 1.9 and Remark 1.10, the condition “$\dim S\leq\frac{1}{2}\dim_{{\mathbb{Q}}}V$” can be removed if ${\mathscr{J}}({\mathbb{V}}^{\prime}_{{\mathbb{Z}}})\rightarrow S$ is not an abelian scheme, with ${\mathbb{V}}^{\prime}_{{\mathbb{Z}}}$ defined above Theorem 1.6. This follows from the comments below Corollary 1.7 (or Theorem 5.2) and is ultimately a consequence of Griffiths’ transversality.
Kerr–Tayou [KT24] also proved Corollary 1.9 and Remark 1.10, with a different method, under the assumption of Theorem 5.2.
1.3. Conjectures on the rationality of the Betti strata, the height inequality, and the volume identity.
Let $S$ be a quasi-projective complex variety, $f\colon X\rightarrow S$ be a smooth family of projective varieties, and $N$ be a relatively ample line bundle on $X$. Let $Z$ be a family of homologically trivial $n$-cocycles in $X/S$ with $n\leq\frac{d+1}{2}$, which is homologically trivial and primitive on each geometric fiber $X_{s}$ in the sense that $c_{1}(N_{s})^{d+2-2n}\cdot Z_{s}=0$ in ${\mathrm{Ch}}^{d+2-n}(X_{s})$. Then we have a normal function $\nu\colon S\rightarrow{\mathscr{J}}^{n}(X/S)_{\mathrm{prim}}$ as defined by (C.5), where ${\mathscr{J}}^{n}(X/S)_{\mathrm{prim}}$ is the intermediate Jacobian associate with the following VHS:
${\mathbb{H}}^{n}:=\ker\left(c_{1}(N)^{d+2-2n}\colon R^{2n-1}f_{*}{\mathbb{Z}}(n){\longrightarrow}R^{2d-2n+3}f_{*}{\mathbb{Z}}(d+2-n)\right).$
We would like to propose the following conjecture on the rationality of the Betti strata.
Conjecture 1.11.
Let $k$ be a subfield of ${\mathbb{C}}$ over which $f$ is defined. Then $S^{\mathrm{Betti}}(t)$ is defined over $k$ for each $t\geq 0$.
In this paper, we prove Conjecture 1.11 for $t=1$ for the Gross–Schoen cycles and Ceresa cycles for any subvariety $S$ of ${\mathscr{M}}_{g}$ by a volume identity (6.4); see Corollary 6.4. We expect this volume identity to be true for general family $f\colon X\rightarrow S$. We point out that in the particular case where $k$ is algebraically closed (for example $k=\overline{\mathbb{Q}}$), some partial results towards Conjecture 1.11 can be deduced from [KOU23], whose proof is by transcendental method.
Now, we assume that $f$ and $N$ are defined over a number field $K$; in general, any family $X/S$ can be embedded into one defined over a number field $K$ by the Lefschetz principle. Assume that the Beilinson–Bloch pairing is well-defined for the cycle $Z_{s}$ for each $s\in S(\overline{\mathbb{Q}})$. This means that $Z_{s}$ can be extended to a cycle ${\mathscr{Z}}_{s}$ on an integral model ${\mathscr{X}}_{s}$ of $X_{s}$ over $O_{K(s)}$ such that the restrictions of ${\mathscr{Z}}_{s}$ on special fibers are homologically trivial. Let $g_{s}$ be the Green’s current on $X_{s}({\mathbb{C}})$ of degree $(n-1,n-1)$ so that ¯$\frac{\partial\bar{\partial}}{\pi i}g_{s}=\delta_{Z_{s}({\mathbb{C}})}$. Then we have an arithmetic cycle $\overline{\mathscr{Z}}_{s}=({\mathscr{Z}}_{s},g_{s})$. The Beilinson–Bloch height of $Z_{s}$ is defined in terms of arithmetic intersection number on ${\mathscr{X}}_{s}$:
(1.8) $\langle{Z_{s},Z_{s}}\rangle_{\mathrm{BB}}=(-1)^{i}\widehat{c}_{1}(\overline{\mathscr{N}}_{s})^{d+1-2n}\overline{\mathscr{Z}}_{s}\cdot\overline{\mathscr{Z}}_{s},$
where $\overline{\mathscr{N}}_{s}$ is a Hermitian line bundle on ${\mathscr{X}}_{s}$ extending $N_{s}$. They even conjectured that this pairing is non-negative and is zero if and only if $Z_{s}$ is torsion in the Chow group. See [Bei87, Blo84, GS91, GS94, Zha22] for details and various connections to the Grothendieck standard conjectures. Besides Gross–Schoen [GS95] for the triple product of curves, we know that the height pairing is unconditionally defined on abelian varieties by Künnemann [Kün01], and the product of a curve and a surface by the second-named author [Zha24].
In the following, we conjecture a height inequality as in our main Theorem 1.1. We say that a function $\mathbf{h}\colon S(\overline{\mathbb{Q}})\rightarrow{\mathbb{R}}$ is a dominant height on an open subvariety $U$ of $S$ over $K$, if there exist a immersion $S\subseteq{\mathbb{P}}^{N}$ defined over $\overline{\mathbb{Q}}$ and positive numbers $\epsilon$ and $c$ such that
(1.9) $\mathbf{h}(s)\geq\epsilon h_{\mathrm{Weil}}(s)-c\qquad\text{ for all }s\in U(\overline{\mathbb{Q}}),$
where $h_{\mathrm{Weil}}$ is the logarithmic Weil height on ${\mathbb{P}}^{N}(\overline{\mathbb{Q}})$. Notice that if this condition is satisfied, then the height inequality (1.9) remains true with $h_{\mathrm{Weil}}$ replaced by $h_{\overline{M}}$ for any integrable adelic line bundle $\overline{M}$ on $S$ (see §$\S$§E.6 for definition) up to changing $\epsilon>0$ and $c>0$.
Write $S^{\mathrm{amp}}=S\setminus S^{\mathrm{Betti}}(1)$, and $S^{\mathrm{slim}}=S^{\mathrm{Betti}}(1)$ with respect to the normal function $\nu$ constructed from $Z$. We conjecture:
Conjecture 1.12.
The Beilinson–Bloch height function $s\mapsto\langle{Z_{s},Z_{s}}\rangle_{\mathrm{BB}}$ is a dominant height on $S^{\mathrm{amp}}$. Moreover, $S^{\mathrm{amp}}$ is the largest open subvariety of $S$ on which $s\mapsto\langle{Z_{s},Z_{s}}\rangle_{\mathrm{BB}}$ is a dominant height.
The second part of Conjecture 1.12 is equivalent to: For any immersion $S\subseteq{\mathbb{P}}^{N}$ defined over $\overline{\mathbb{Q}}$, there exists a Zariski dense sequence $s_{i}\in S^{\mathrm{slim}}(\bar{K})$ such that
$\lim_{i}h_{\mathrm{Weil}}(s_{i})=\infty,\qquad\lim_{i}\frac{\langle{Z_{s_{i}},Z_{s_{i}}}\rangle_{\mathrm{BB}}}{h_{\mathrm{Weil}}(s_{i})}=0.$
The proof of Theorem 1.1 executed in §$\S$6.3 in fact proves the first part of Conjecture 1.12 in the case of Gross–Schoen and Ceresa cycles for any subvariety of ${\mathscr{M}}_{g}$. A key ingredient is the volume identity (6.4). The same argument can be used to show that the following conjectured volume identity (1.10) implies the full Conjecture 1.12 with the help of [YZ21, Thm. 6.3.6]. To check whether $S^{\mathrm{amp}}$ is non-empty, one can use (1.6) and Theorem 1.6.
In [Blo89], Bloch constructed a biextension line bundle $L$ on $S$ defined over $K$. At each closed point $s\in S(\overline{\mathbb{Q}})$ where the Beilinson–Bloch height is well-defined, this line bundle has an adelic structure $\overline{L}_{s}$ over $K(s)$ whose degree gives the Beilinson–Bloch height (1.8). The archimedean part of this adelic structure has been studied by Hain [Hai90].
Conjecture 1.13.
The line bundle $L$ has a natural extension to a nef adelic line bundle $\overline{L}$ on $S$ in the sense of [YZ21], i.e. the image of $\overline{L}$ under the composite of (E.6) and (E.5) is $L$, such that its restriction to each $s\in S(\overline{\mathbb{Q}})$ where the Beilinson–Bloch height is well-defined is given by $\overline{L}_{s}$.
By the arithmetic Hilbert–Samuel formula [YZ21, Thm. 5.2.2] (which we recall as Theorem E.8.(i)), one consequence of this conjecture is the following volume identity for the generic fiber $\widetilde{L}$ of $\overline{L}$ (see (E.6) for the definition): For any archimedean place $v$ of $K$ and any subvariety $S^{\prime}$ of $S_{v,{\mathbb{C}}}$, we have
(1.10) $\widehat{\mathrm{vol}}(\widetilde{L}|_{S^{\prime}})=\int_{S^{\prime}({\mathbb{C}})}c_{1}(\bar{L}_{v})^{\dim S}.$
For the Gross–Schoen and Ceresa cycles for any subvariety of ${\mathscr{M}}_{g}$, we prove the existence of the extension in §$\S$6.1 and the volume identity as Theorem 6.3, but not the nefness of the adelic line bundle in question.
@@ 1.4. Plan and Ingredients of proofs
The arithmetic part requires the construction of a suitable adelic line bundle $\overline{{\mathscr{L}}}$ over the base, a theory initiated by the second-named author [Zha95] and developed in further joint work with Yuan [YZ21], such that the height function $h_{\overline{{\mathscr{L}}}}$ is the Beilinson–Bloch height. Then, the desired height inequalities will follow from suitable bigness properties of $\overline{{\mathscr{L}}}$. In the context of Theorem 1.1 and Corollary 1.2, by the second-named author’s [Zha10, Thm. 2.5.5] the desired $\overline{{\mathscr{L}}}$ can be obtained from a suitable Deligne pairing. This $\overline{{\mathscr{L}}}$ has been constructed by Yuan [Yua21, §$\S$3.3.4] is slightly different steps. The required bigness property is the bigness of the generic fiber $\widetilde{{\mathscr{L}}}$ of $\overline{{\mathscr{L}}}$. We furthermore prove the volume identity as Theorem 6.3, which implies that the bigness of $\widetilde{{\mathscr{L}}}$ will follow from the non-vanishing of $c_{1}(\overline{{\mathscr{L}}})^{\wedge\dim S}$ with $c_{1}(\overline{{\mathscr{L}}})$ the curvature form. The arithmetic part of this paper is confined to the Gross-Shoen and the Ceresa cycles.
@@
The geometric part studies the (admissible) normal function $\nu$ associated with $Z$. The main results are the Zariski closedness of the Betti strata and the formula for the Betti rank $r(\nu)$. This part has been explained in §$\S$1.2. The proofs of Theorem 1.4 and Theorem 1.6 are simultaneous and follow the guideline of the first-named author’s [Gao20a] on the generic rank of the Betti map for abelian schemes. A core of our proof is Ax–Schanuel for VMHS independently proved by Chiu [Chi24] and Gao–Klingler [GK24], which will be used multiple times.
@@
Hain’s works on the Hodge-theoretic computation of the Archimedean local height pairing and the Betti form are the key ingredients to bridging the geometric and arithmetic parts. Indeed, Hain in [Hai90] proved that the archimedean local height pairing could be computed using the metrized biextension line bundle on $S$, and in [HR04] computed (joint with Reed), the curvature form of the metrized biextension line bundle. In our situation, we work with the height pairing of $Z_{s}$ with itself for any $s\in S(\overline{\mathbb{Q}})$, and the metrized biextension line bundle in question is the pullback of the metrized tautological line bundle under $\nu$. Hain in [Hai13] proved that its curvature form $\beta_{\nu}$, called the Betti form, is a semi-positive $(1,1)$-form. We show in Corollary D.7 that $\beta_{\nu}^{\wedge\dim S}\not\equiv 0$ if and only if $r(\nu)=\dim S$. Finally, back to the situation of Theorem 1.1, Corollary 1.2, and Theorem 1.3. Hain’s formula [Hai90, Prop. 3.3.12] and the second-named author’s [Zha10, Thm. 2.5.5] together imply that the curvature $c_{1}(\overline{{\mathscr{L}}})$ is precisely the Betti form. This finishes the bridge between the geometric part and the arithmetic part. Again, the bridge between the geometric part and the arithmetic part works in much larger generality, provided that the arithmetic part is solved.
1.5. Organization of the paper
We will work on the geometric part of our proof in §$\S$25. More precisely, we will define the period map associated with a normal function in §$\S$2 and explain how to see the Betti foliation in this language, prove the formula to compute the Betti rank in §$\S$3 assumption a finiteness result à la Ullmo which will be proved in §$\S$4, and then explain two applications of this formula in geometric situations (including the Gross–Schoen and Ceresa normal functions as particular cases) in §$\S$5. Then, in §$\S$6, we will explain the arithmetic part on how to construct the desired adelic line bundle, prove the volume identity, and prove our main theorem by combining the results of both the arithmetic and the geometric part.
For readers’ convenience, we include several appendices, for which we do not claim originality, to make our paper as self-contained as possible.